OpenStudy (anonymous):

Use the method of substitution to solve the system of linear equations. 5x+5y=5/3 y=3x-1 the solution is an ordered pair

OpenStudy (anonymous):

5x+5(3x-1) = 5/3 distribute and collect terms, and solve for x x=(1/3) y=3x-1 y=3(1/3)-1 y=0 solution is (1/3,0)

OpenStudy (anonymous):

thank you so much you are a lifesaver